### v2

# COSMOS: Cross-Modality Self-Distillation for Vision Language Pre-training

Abstract
Vision-Language Models (VLMs) trained with contrastive loss have achieved significant advancements in various vision and language tasks. However, the global nature of the contrastive loss makes VLMs focus predominantly on foreground objects, neglecting other crucial information in the image, which limits their effectiveness in downstream tasks. To address these challenges, we propose COSMOS: CrOSs-MOdality Self-distillation for vision-language pre-training that integrates a novel text-cropping strategy and cross-attention module into a self-supervised learning framework. We create global and local views of images and texts (i.e., multi-modal augmentations), which are essential for self-distillation in VLMs. We further introduce a cross-attention module, enabling COSMOS to learn comprehensive cross-modal representations optimized via a cross-modality self-distillation loss. COSMOS consistently outperforms previous strong baselines on various zero-shot downstream tasks, including retrieval, classification, and semantic segmentation. Additionally, it surpasses CLIP-based models trained on larger datasets in visual perception and contextual understanding tasks. Code is available at https://github.com/ExplainableML/cosmos.
1 Introduction
CLIP [63] uses an instance-level contrastive loss [60] to learn text and image representations, where the text gets embedded close to its corresponding image. The contrastive loss function employed to train CLIP is inherently global, as it matches the entire image to the text. This leads to the dominant objects in the image suppressing the recognition of other smaller objects [12, 64, 90, 44], which results in poor performance for dense prediction tasks, such as semantic segmentation [57], and failure to distinguish similar-looking images with different visual patterns [74]. CLIP also demonstrates a lack of contextual understanding, as its text encoder perceives a caption as a bag-of-words, ignoring the sequence of words [73, 87, 33].
Figure 1: We randomly crop the image and randomly select captions (1-5 sentences) to build global and local crops (top). In our cross attention module, two modalities are conditioned on each other to create attention maps (normalized to 0-1, bottom).
Previous works [45, 56, 57, 21, 65] have tackled these limitations by masking the text or the image [21, 45, 65] and explicitly optimizing local-to-global image representation. For instance, SLIP [56] and SILC [57] improved the visual representations by integrating self-supervision into the CLIP image encoder. However, we argue that enhancing only the image encoder leads to sub-optimal results, since this does not fully exploit the multi-modality of the data. Therefore, we propose COSMOS, to learn CrOSs-MOdality embeddings through Self-distillation. Unlike previous self-supervised methods [56, 57, 21], COSMOS incorporates a unique text-cropping strategy and a cross-attention module to learn multi-modal representations for various downstream tasks.
Our text-cropping strategy (Fig. 1, top) is inspired by the multi-crop augmentation for images [7], which optimizes local-to-global image correspondence during training. We randomly sample from the long synthetic caption datasets [91, 70] to construct sentences of varying length. By applying multi-crop augmentation to both images and text, we enhance the visual and text encoders simultaneously within a self-distillation framework [7], aligning the representation of the student VLM with that of the teacher.
Our cross-attention module, integrated into the student, allows the model to attend to the other modality’s representation, ensuring that cross-modality information is captured during training. Inside the module, image patch tokens are conditioned on texts, or word tokens are conditioned on images, to learn visual and textual grounding. As illustrated in Fig. 1 (bottom), our cross-attention module focuses on different objects from the entire image, which are highlighted in both images and captions based on the attention weights. This demonstrates that COSMOS is able to effectively localize relevant information in images and captions.
We demonstrate the adaptability of COSMOS by pre-training it on 4 datasets varying from 3M to 30M pairs. To evaluate our method, we perform zero-shot experiments on 2 retrieval, 11 classification, and 6 segmentation benchmarks, as well as, the visual perception [74] and contextual understanding tasks [33, 32]. The extensive experiments on these downstream tasks reveal that our method consistently outperforms the previous CLIP-based methods, under the same pre-training data budget. Moreover, COSMOS pre-trained on only 12 million examples outperforms the OpenCLIP [13] models trained on billions of image-text pairs, in both image-to-text and text-to-image retrieval tasks.
Our contributions are summarized as follows: (1) We design a text-cropping strategy which promotes local-to-global text representation learning, fully leveraging multi-modality in the self-distillation of VLMs. (2) We demonstrate that cross-modality embeddings are effective to simultaneously self-distill both image and text encoders. (3) COSMOS reveals fine-grained perception of images and compositional understanding of language, outperforming baselines trained on larger datasets.
2 Related Work
Figure 2: An overview of COSMOS. Left: Our VLM pre-training mechanism is based on the student-teacher framework with contrastive loss ($\mathcal{L}_{\text{CLIP}}$) for multi-modal alignment, and cross-modality self-distillation loss ($\mathcal{L}_{\text{COSMOS}}$) for fine-grained representation learning. Right: The architecture of the student and teacher model with cross-attention modules that extract cross-modality information from the student.
Self-Supervised Learning (SSL) learns discriminative features from unlabeled data [30]. Initially, models like BERT [20] and GPT [62] were designed to predict masked words or the next word in a sequence, eliminating the need for additional labels. Similarly, vision models benefit from self-supervised techniques, either with contrastive learning or masked image modeling. Based on contrastive learning, BYOL [29] and DINO [7] trained a student network to produce image embeddings that closely resemble those of a teacher network, with the teacher being an exponential moving average of the student. On the other hand, masked image modeling divides the input image into visual tokens and predicts a randomly masked subset of them. For example, MAE [31] directly predicted the original pixel from the masked patches, and BEiT [3] employed a pre-trained tokenizer as its target.
Contrastive learning in vision-language pre-training has achieved great success in various downstream tasks, demonstrating strong zero-shot transfer abilities [63, 35, 83, 26, 28, 42, 27]. As these models are trained to learn visual representations aligned with the language representation, most Multi-modal Large Language Models (MLLMs) [1, 43, 50, 49, 94] adopt these vision encoders to process the image input. However, instance-level contrastive learning suffers from feature suppression, where the model learns only the dominant features in the data while neglecting other valuable features [12, 64, 90, 44, 71, 81, 47, 74]. In other words, the model creates so-called simple shortcut features and decision rules that do not consider all available distinguishing features. Some works [10, 70, 75, 89, 59, 91, 78] have employed datasets equipped with long synthetic captions generated by MLLMs. These datasets help mitigate feature suppression by learning different features described in the captions and avoiding overfitting to a single caption. Referring to Zheng et al. [91], we employ image datasets with long synthetic captions. Instead of simply training our model with these additional captions, we create text augmentation using the synthetic captions, and introduce the cross-attention module for self-distillation.
Following the success of SSL, recent works [45, 56, 57, 21, 65, 37] apply SSL objectives to contrastive vision-language pre-training. These approaches also tackle the feature suppression limitation, where the new objective reduces the inductive bias to prevent the creation of shortcut features. For example, SLIP [56] combined the SimCLR [11] loss with the contrastive loss, resulting in better transferability compared to the original CLIP. Similarly, SILC [57] explicitly matched local crops to global crops during training inspired by DINO [7], while MaskCLIP [21] learned more fine-grained features by applying masked modeling to both images and text. Similar to SLIP [56] and SILC [57], we also employ self-supervised learning for vision-language pre-training. While SLIP [56] and SILC [57] only improve the image representation without improving the text representation, we propose a text-cropping technique that enables us to improve the text and image representations simultaneously with our novel cross-modality self-distillation loss.
3 COSMOS Model
We describe our augmentation technique, the key component in self-distillation, in Sec. 3.1, COSMOS’s architecture in Sec. 3.2, and the training objective in Sec. 3.3.
3.1 Text and Image Augmentations
Text Cropping Strategy. To perform self-supervised learning based on text, we propose a text-cropping strategy utilizing long caption datasets [91, 70]. These datasets consist of images paired with detailed synthetic captions generated by MLLMs. Inspired by the multi-crop strategy for images [7], we define global and local concepts for text captions. The global crop of a caption comprises one to five sentences randomly sampled from the long synthetic caption, while the local crop of a caption consists of a single sentence, also randomly sampled from the long caption. The number of sentences in the global captions is randomly determined at each training iteration. This adaptation ensures that the global caption generally describes larger areas of the original image, e.g. the fox with vibrant red fur is the main subject of the image, in Fig. 1 top, while the local caption focuses on relatively smaller regions, e.g. the snow covers the ground and the trees.
Image Cropping Strategy. Following [7], we adopt the standard multi-crop strategy [6] for images as illustrated in Fig. 1 top, utilizing two global views and several local views. Both local and global views are processed through the student encoder, while only the global views are passed through the teacher encoder, thereby promoting local-to-global correspondences. It is noteworthy that the image cropping and text cropping strategies are executed independently, meaning their global and local crops might not contain the same information. For instance, in Fig. 1 top, the local image crop displays the fox’s tail, while the local text crop describes the image’s background. This configuration is the standard setting of COSMOS, unless otherwise stated.
3.2 Model Architecture
The overall architecture of COSMOS is depicted in Fig. 2. Both the student model $\textbf{G}_{\theta_{s}}$ and the teacher model $\textbf{G}_{\theta_{t}}$ are Vision-Language Models (VLMs), each comprising an image encoder $\textbf{I}_{\theta}$ and a text encoder $\textbf{T}_{\theta}$, such that $\textbf{G}_{\theta_{*}}=\{\textbf{I}_{\theta_{*}},\textbf{T}_{\theta_{*}}\}$, where $*\in\{s,t\}$. The student and teacher models share the same architecture and are initialized with identical weights. Additionally, the student model includes a cross-attention module $\textbf{C}_{\theta}=\{\textbf{C}^{I}_{\theta},\textbf{C}^{T}_{\theta}\}$, which extracts cross-modal embeddings.
As shown on the left side of Fig. 2, two random multi-modal transformations are applied to an image-text pair $p$, resulting in $\tilde{p}_{s}$ and $\tilde{p}_{t}$. The student model $\textbf{G}_{\theta_{s}}$ processes $\tilde{p}_{s}$, followed by the cross-attention module and normalization operation to produce the output $f_{s}$. Similarly, the teacher model $\textbf{G}_{\theta_{t}}$ processes $\tilde{p}_{t}$ to generate the output $f_{t}$, which serves as the target. The student model, along with the cross-attention module, is optimized by matching its output $f_{s}$ with the teacher’s output $f_{t}$ ($\mathcal{L}_{\text{COSMOS}}$), while standard contrastive loss is calculated within the student ($\mathcal{L}_{\text{CLIP}}$). A stop gradient operation is applied to the teacher model to ensure that gradients propagate only through the student. The teacher parameters $\theta_{t}$ are updated at each iteration using the exponential moving average (EMA) of the student parameters $\theta_{s}$ (i.e., $\theta_{t}=\lambda\theta_{t}+(1-\lambda)\theta_{s}$ with the parameter $\lambda$ defining the update schedule). This learning mechanism, combined with multi-modal augmentations, enables COSMOS to learn the correspondence between local features from the student and global patterns from the teacher.
[…]
3.3 Training Objective
[…]
$\mathcal{L}_{\text{nce}}(x,y)=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp\left({\langle x_{i},y_{i}\rangle/\tau}\right)}{\sum_{j=1}^{B}\exp\left({\langle x_{j},y_{i}\rangle/\tau}\right)}$ (3)
[…]
$\mathcal{L}_{\text{sym-nce}}(x,y)=\frac{1}{2}(\mathcal{L}_{\text{nce}}(x,y)+\mathcal{L}_{\text{nce}}(y,x)).$ (4)
The total contrastive loss $\mathcal{L}_{\text{CLIP}}$ is calculated between the image and text embeddings of the student as follows.
$\mathcal{L}_{\text{CLIP}}=\mathcal{L}_{\text{sym-nce}}(\text{CLS}_{s},\text{EOT}_{s}).$ (5)
[…]
$\mathcal{L}_{\text{COSMOS}}=\frac{1}{4}(\mathcal{L}_{\text{sym-nce}}(h_{I},\text{CLS}_{t})+\mathcal{L}_{\text{sym-nce}}(h_{I},\text{EOT}_{t})\\ +\mathcal{L}_{\text{sym-nce}}(h_{T},\text{CLS}_{t})+\mathcal{L}_{\text{sym-nce}}(h_{T},\text{EOT}_{t}))$ (6)
[…]
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CLIP}}+\mathcal{L}_{\text{COSMOS}}.$ (7)
As mentioned in Sec. 3.1, all crops are passed through the student while only global crops are given to the teacher. Thus, $\mathcal{L}_{\text{CLIP}}$ is calculated between all crops within the student, optimizing the model to construct the overall structure of image and text representations by aligning the embeddings through contrastive learning. On the other hand, $\mathcal{L}_{\text{COSMOS}}$ is obtained between global crops given to the teacher and all crops given to the student, encouraging the model to learn fine-grained representations by predicting the rich global context of the teacher from the local feature of the student (the detailed pseudocode of the training objective is in Supplementary Material Algorithm 1).
Flickr30K MSCOCO
Image-to-text Text-to-image Image-to-text Text-to-image
Data Method R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CC3M MLLM-A [51] 63.5 86.6 91.7 49.3 74.8 83.1 35.9 62.4 73.9 26.5 51.1 62.7
CLIP [63] 68.4 88.9 93.2 52.1 78.4 85.5 40.2 67.0 77.5 27.2 53.5 65.0
SigLIP [88] 69.8 90.4 94.5 53.1 78.5 86.1 40.1 67.5 78.1 28.3 54.4 65.9
DreamLIP [91] 69.2 91.5 95.9 53.5 78.3 85.5 40.8 68.5 79.4 29.8 55.4 66.3
COSMOS 84.1 96.2 98.9 68.6 88.7 93.4 53.1 78.3 86.7 40.1 66.8 77.1
CC12M SLIP [56] 62.5 87.2 92.1 46.6 73.3 80.9 37.6 64.9 75.5 26.8 51.4 62.7
LaCLIP [23] 63.9 86.5 92.6 51.6 78.8 86.2 38.0 64.8 75.0 26.5 51.2 62.6
LaSF-CLIP [65] 71.8 91.9 95.2 59.9 84.2 90.9 44.3 71.3 80.2 31.4 57.3 68.1
CLIP [63] 81.8 96.5 98.1 66.9 87.7 92.7 56.2 80.4 88.0 39.8 66.4 76.6
SigLIP [88] 82.5 96.1 97.8 66.7 88.1 92.8 55.4 79.7 87.8 40.4 67.0 76.6
DreamLIP [91] 84.1 97.8 99.0 68.3 89.4 93.4 54.0 78.2 86.4 40.7 66.5 76.4
COSMOS 91.4 98.2 98.8 76.2 92.9 96.3 64.2 85.6 91.6 48.9 74.3 82.9
YFCC15M SLIP [56] 58.6 85.1 91.7 41.3 68.7 78.6 33.4 59.8 70.6 21.5 44.4 56.3
MaskCLIP [21] 70.1 90.3 95.3 45.6 73.4 82.1 41.4 67.9 77.5 25.5 49.7 61.3
SF-CLIP [65] 68.7 90.4 94.8 46.2 73.2 82.7 41.8 68.3 78.4 25.5 50.2 61.3
CLIP [63] 89.1 97.6 98.8 72.4 90.7 94.9 61.0 85.0 90.9 44.7 71.2 80.3
SigLIP [88] 89.9 97.5 98.9 73.6 92.1 95.5 62.5 85.3 91.8 46.5 72.8 81.9
DreamLIP [91] 87.3 98.1 99.0 70.0 89.2 93.4 57.0 80.9 88.4 42.4 68.5 78.2
COSMOS 92.6 99.1 99.7 79.6 94.2 96.8 67.5 88.2 93.5 50.9 76.0 84.2
Merged-30M CLIP [63] 90.5 98.7 99.4 75.9 93.5 96.7 63.2 85.6 91.7 48.2 73.9 82.4
SigLIP [88] 90.5 98.3 99.4 76.7 93.2 96.1 64.1 86.2 91.9 49.0 74.5 83.2
DreamLIP [91] 89.9 99.0 99.6 73.3 91.9 95.7 62.3 84.5 90.3 44.9 69.9 79.2
COSMOS 92.9 99.4 99.9 80.3 95.3 97.6 68.0 87.8 92.5 52.5 77.2 84.9
LAION-400M OpenCLIP [13] 86.0 97.3 98.9 68.1 88.6 93.3 56.5 80.4 87.3 37.9 63.2 73.3
DataComp-1B OpenCLIP [13] 84.5 96.8 98.4 67.9 88.0 93.0 58.2 80.8 88.5 39.8 65.5 75.0
LAION-2B OpenCLIP [13] 87.6 97.7 99.5 71.9 90.5 94.8 59.3 82.5 89.0 41.7 67.1 76.5
2.5B MetaCLIP [80] 85.9 97.3 98.9 70.5 90.7 94.6 59.4 80.6 87.9 41.4 67.2 77.0
Llip [41] 90.1 98.5 99.6 75.1 92.8 96.2 63.4 84.3 90.3 45.6 70.8 79.7
Table 1: Zero-shot image-text retrieval results in terms of R@1, R@5, and R@10 on the Flickr30K [84] and MSCOCO [48] datasets. The vision encoder architecture is ViT-B/16. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
Simple Loss Formulation. Previous works [45, 57, 21, 65] often rescale the CLIP loss when combining it with distillation objectives, to ensure that the new loss does not disrupt the original learning mechanism of CLIP. Simply adding self-distillation loss without scaling the CLIP loss often underperforms the original model. To overcome this limitation, one must grid-search for the optimal loss scale factor. In our method, both CLIP and the COSMOS loss are updated at the same scale, eliminating the need for additional hyperparameter tuning to adjust the ratio of the two losses.
4 Experiments
Data Model
Food-101
CIFAR-10
CIFAR-100
SUN397
Cars
Aircraft
DTD
Pets
Caltech-101
Flowers
ImageNet
Average
CC3M LaCLIP [23] 14.2 57.1 27.5 35.1 1.6 1.6 16.6 15.6 52.7 14.7 21.5 23.5
MLLM-A [51] 18.7 58.4 32.4 43.8 3.9 1.5 20.2 32.1 63.5 17.5 25.0 28.8
CLIP [63] 17.9 75.0 40.8 43.1 2.6 1.0 15.3 22.1 68.9 12.6 23.9 29.4
SigLIP [88] 18.4 76.4 41.9 46.9 3.0 1.4 17.6 20.6 70.4 10.8 25.9 30.3
DreamLIP [91] 23.6 75.7 44.2 46.1 3.6 1.6 18.7 27.4 73.6 18.5 31.6 33.1
COSMOS 31.2 79.1 51.6 54.2 5.2 2.3 27.4 37.5 78.1 25.6 37.1 39.0
CC12M LaCLIP [23] 60.7 75.1 43.9 57.0 36.3 5.6 31.0 72.4 83.3 39.9 48.4 50.3
LaSLIP [23] 62.9 82.0 50.2 59.6 32.2 4.4 30.1 70.6 82.4 37.4 49.7 51.0
MLLM-A [51] 60.9 83.0 55.4 59.4 24.1 3.2 30.7 64.8 79.3 36.0 47.5 49.5
CLIP [63] 38.4 90.4 58.6 57.7 12.2 2.0 24.6 40.9 80.8 12.1 39.1 41.5
SigLIP [88] 42.3 88.9 59.6 57.6 11.2 2.3 24.1 38.4 80.0 13.5 38.9 41.5
DreamLIP [91] 59.2 87.7 63.8 54.5 29.9 5.1 29.5 60.8 82.5 32.5 50.2 50.5
COSMOS 63.1 92.8 70.8 64.2 30.4 3.4 36.5 56.9 85.5 36.3 51.4 53.8
YFCC15M SLIP [56] 59.9 82.6 48.6 - 5.4 5.6 26.6 31.4 70.9 56.3 42.8 -
MaskCLIP [21] 64.9 80.2 57.5 - 6.7 6.1 27.9 34.3 72.0 57.0 44.5 -
SF-CLIP [65] 65.0 85.0 53.6 - 8.2 11.0 35.2 38.1 72.2 59.5 45.2 -
CLIP [63] 43.5 88.0 56.8 55.8 9.3 4.7 26.0 27.9 76.7 37.0 41.7 42.5
SigLIP [88] 45.7 87.8 58.7 58.6 9.9 4.7 28.1 32.3 79.1 36.4 44.4 44.2
DreamLIP [91] 45.0 88.9 62.8 57.5 9.4 5.8 32.0 33.6 81.5 44.4 48.2 46.3
COSMOS 55.8 90.4 65.6 59.7 16.8 5.8 33.7 39.3 83.1 46.9 52.4 50.0
Merged-30M CLIP [63] 61.3 92.2 66.9 62.2 19.3 5.7 30.9 49.3 83.7 43.4 50.0 51.4
SigLIP [88] 64.2 91.0 67.6 64.0 22.0 5.7 33.5 53.3 84.3 43.6 51.0 52.7
DreamLIP [91] 76.4 92.7 71.7 64.7 23.4 8.3 33.9 66.3 87.6 59.1 58.4 58.4
COSMOS 73.9 93.4 72.3 67.8 31.2 5.9 37.3 62.9 88.0 54.6 57.6 58.6
LAION-400M OpenCLIP [13] 86.1 91.7 71.4 69.0 83.7 17.9 50.7 89.3 90.1 69.3 67.1 71.5
DataComp-1B OpenCLIP [13] 90.5 96.3 82.2 70.0 88.9 29.8 57.7 92.8 91.8 75.7 73.5 77.2
LAION-2B OpenCLIP [13] 86.6 94.9 76.9 70.5 88.5 26.8 56.5 90.2 90.9 71.4 70.2 74.9
2.5B MetaCLIP [80] 88.3 95.7 79.0 68.5 82.9 30.3 62.1 91.7 93.3 73.9 72.1 76.2
Llip [41] 89.0 95.7 81.4 70.9 88.2 41.5 63.7 93.5 94.7 74.9 75.3 79.0
Table 2: Zero-shot classification results in terms of top-1 accuracy on ImageNet [18] and Food101 [4], CIFAR-10 [39], CIFAR-100 [39], SUN397 [79], Stanford Cars [38], FGVCAircraft [54], DTD [14], Oxford Pets [61], Caltech101 [24], Flowers102 [58], and ImageNet [18]. The vision encoder architecture is ViT-B/16. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
Pre-training Datasets. We pre-train COSMOS on Conceptual Caption 3M (CC3M) and 12M (CC12M) [68] as well as YFCC15M [16], a subset of YFCC100M [72]. The combination of these three datasets (CC3M, CC12M, and YFCC15M) results in a merged dataset of 30M pairs, referred to as Merged-30M, following DreamLIP [91]. As mentioned in Sec. 3.1, we utilize synthetic long captions to obtain multi-modal augmentation. For CC3M, CC12M and YFCC15M, we use the synthetic captions generated by DreamLIP [91]. The new captions were obtained using InstructBLIP [17], LLaVA-1.5 [49], and ShareGPT4V [10].
Implementation Details. COSMOS is built on the OpenCLIP [13] code repository, where we primarily adopt the default settings of their implementation. We use ViT-B/16 as the vision encoder designed for $224\times 224$ image sizes, while our CLIP text encoder has the sequence length of 77. Each encoder consists of 12 transformer layers, and the embedding size is fixed at 512. We train COSMOS on A100 GPUs for 32 epochs, using a learning rate of $5\times 10^{-4}$ and a batch size of 1,024 for CC3M, or 4,096 for CC12M, YFCC15M and Merged-30M. Following DINO [7], we crop two global image views at scales between 0.4 and 1.0, resizing them to $224\times 224$ pixels, and six local image views at scales between 0.05 and 0.4, resized to $96\times 96$ pixels. Similarly, the two global text views consist of one to five sentences randomly sampled from the synthetic captions, while the six local text views consist of a single sentence sampled from the same caption set. The teacher momentum $\lambda$ is empirically set to 0.999 for CC3M and 0.99 for other datasets (more details in Sec. C.1).
4.1 Zero-Shot Image-Text Retrieval
As shown in Tab. 1, COSMOS achieves 53.1% and 40.1% R@1 score for image-to-text (I2T) and text-to-image (T2I) retrieval on MSCOCO [48]. These results are 12.3% and 10.3% higher than DreamLIP [91] (40.8% and 29.8%, respectively). When we switch to a larger pre-training set (CC12M), COSMOS further improves its performance, reaching 64.2% and 48.9% R@1 scores, outperforming CLIP [63] by 8.0% and 9.1%, respectively. When we employ YFCC15M [16] as the pre-training dataset, COSMOS attains a performance of 67.5% and 50.9% in terms of I2T and T2I R@1 scores. The highest performance of 68.0% and 52.5% on the MSCOCO [48] dataset is obtained by COSMOS, when Merged-30M is used for pre-training. COSMOS also achieves the best performance for image-text retrieval on the Flickr30k [84] dataset, achieving an R@1 score of 92.9% and 80.3% in terms of I2T and T2I retrieval when Merged-30M is used for pre-training, while SigLIP [88] attains only 89.9% and 73.3%, respectively. Overall, COSMOS outperforms all baselines on both datasets (Flickr30k [84] and MSCOCO [48]) in the image-text retrieval task.
We include other baselines for state-of-the-art comparison, including CLIP-based models (e.g., OpenCLIP [13], MetaCLIP [80], and Llip [41]) trained on larger datasets and models based on text augmentation (e.g., LaCLIP [23] and MLLM-A [51]) and self-supervised learning (e.g., SLIP [56], MaskCLIP [21], and SF-CLIP [65]). It is noteworthy that COSMOS trained on Merged-30M even outperforms Llip [23] trained on 2.5B image-text pairs (68.0% vs 63.4% and 52.5% vs 45.6% in terms of R@1 I2T and T2I retrieval on MSCOCO). COSMOS’s improvement stems from the cross-modality distillation, since the distillation signal is passed through both encoders encouraging multi-modal alignment.
Method Data Size With background category Without background category Avg.
VOC21 Context60 COCO-Obj. VOC20 City. Context59 ADE COCO-Stf.
OpenCLIP [13] 400M 18.4 9.1 8.7 43.8 6.9 9.8 2.8 6.9 13.3
OpenCLIP [13] 1B 22.5 11.9 11.4 51.3 8.5 12.6 4.2 9.2 16.5
OpenCLIP [13] 2B 19.8 9.6 9.9 51.3 6.2 10.2 3.5 7.2 14.7
COSMOS 30M 29.6 12.5 15.2 53.6 13.9 15.7 8.5 10.7 20.0
CLIP [63] 400M 18.8 9.9 8.1 49.4 6.5 11.1 3.1 5.7 14.1
MaskCLIP [93] 400M 43.4 23.2 20.6 74.9 24.9 26.4 11.9 16.7 30.3
SCLIP [77] 400M 59.1 30.4 30.5 80.4 32.2 34.2 16.1 22.4 38.2
SILC-C[57] 10B - - - 77.5 26.9 31.6 19.3 20.8 -
COSMOS w/ SCLIP 30M 52.8 31.2 31.3 77.7 34.7 33.7 17.7 23.2 37.8
Table 3: Zero-shot semantic segmentation results in terms of mIoU (mean Intersection over Union) on the PASCAL VOC 2012 [22], PASCAL Context [55], Cityscapes [15], ADE20k [92], COCO-Stuff, and COCO-Object [5] datasets with and without background category. The vision encoder architecture is ViT-B/16. We also report the performance of our method enhanced with the SCLIP mechanism [77]. The best results are highlighted in bold. : Results copied from their work. : Results obtained using their official pre-trained weights.
4.2 Zero-Shot Image Classification
We observe in Tab. 2 that COSMOS outperforms the baselines CLIP [63], SigLIP [88], DreamLIP [91], MaskCLIP [21], and SF-CLIP [65] on most of the evaluation datasets with the same pre-training dataset. For instance, COSMOS pre-trained on CC3M reaches much higher accuracy of 37.1% on ImageNet compared to CLIP (23.9%) and SigLIP (25.9%). As expected, the classification task greatly benefits from the pre-training dataset size, i.e., COSMOS pre-trained on CC12M achieves 51.4% on ImageNet while on Merged-30M, it reaches 57.6%.
Additionally, COSMOS surpasses all baselines regardless of the pre-training set, in terms of average accuracy computed over the 11 datasets. For instance, when pre-trained on YFCC15M, COSMOS achieves an average accuracy of 50.0%, while SigLIP [88] and DreamLIP [91] reach 44.2% and 46.3%, respectively. It is worth noting that on two fine-grained datasets, i.e., Cars and SUN397, the accuracy improvements of COSMOS trained on Merged-30M is significant (67.8% vs 64.7% on SUN397, 31.2% vs 23.4% on Cars compared to DreamLIP).
Models trained on billions of samples, e.g., OpenCLIP [13], MetaCLIP [80], and Llip [41] achieve higher accuracy, implying that the larger the number of images the higher the accuracy for the classification task, since the model is required to learn the global information.
4.3 Zero-Shot Semantic Segmentation
As shown in Tab. 3 top, COSMOS trained on only 30M samples outperforms the best OpenCLIP model [13] trained on 1B samples in all benchmarks. In the datasets with a background category, COSMOS (29.6%, 12.5%, and 15.2%, respectively) surpasses the performance of OpenCLIP (1B) model (22.5%, 11.9%, and 11.4%, respectively) by a large margin. Remarkably, COSMOS almost doubles the performance on the fine-grained Cityscapes dataset [15] and achieves 20.0% average mIoU over all segmentation benchmarks compared to 16.5% obtained by the OpenCLIP model (1B). While CLIP-based models often suffer from feature suppression, COSMOS alleviates this issue through cross-modality embeddings and local-to-global matching.
In addition, we apply SCLIP [77] to our method where the weight of the last transformer block is adapted to construct the Correlative Self-Attention (CSA) block. Wang et al. [77] claimed that the CSA block maximizes the localization of patch embeddings without extra training. Although the CSA mechanism of SCLIP is tailored for the original CLIP model, it also improves the segmentation results of COSMOS. As a result, our COSMOS, trained on 30M samples, achieves an average mIoU of 37.8%, which is comparable to the 38.2% achieved by SCLIP [77], which was trained on 400M samples. Notably, COSMOS outperforms SILC [57] in several benchmarks despite a significantly smaller training budget (30M vs 10B samples). For instance, our model (vs SILC) achieves 34.7% and 23.2% (vs 26.9% and 20.8%) on the Cityscapes and COCO-stuff datasets. These results suggest that COSMOS can effectively generate fine-grained multi-modal representations with a much smaller dataset size. In Sec. E.8, we present additional segmentation results comparing COSMOS to baselines trained on Merged-30M.
Method Data SugarCrepe SVO MMVP
Size Re. Sw. Add Avg Sub. Verb Obj. Avg
CLIP [63] 30M 76.2 68.5 79.5 76.5 52.2 36.4 62.0 18.5
SigLIP [88] 30M 76.6 67.8 77.9 76.0 52.3 36.3 61.5 20.7
DreamLIP [91] 30M 82.7 76.5 82.1 81.8 57.1 42.2 66.5 25.2
COSMOS 30M 87.3 79.1 88.1 86.6 59.0 45.1 68.5 25.9
400M 81.6 65.2 82.5 80.0 58.5 37.1 66.0 25.9
OpenCLIP [13] 1B 82.7 64.6 86.6 81.9 55.5 35.3 62.9 25.9
2B 84.7 67.0 87.4 83.5 58.5 39.5 65.3 21.5
Table 4: Robustness assessment. We evaluate COSMOS on SugarCrepe [33] (Replace = Re., Swap = Sw., Add), SVO [32] (Subject = Sub., Verb, Object = Obj.), and MMVP-VLM [74].
4.4 Visual Perception & Contextual Understanding
We evaluate COSMOS on the visual perception and contextual understanding tasks [74, 33, 32] to demonstrate the robustness of our representations. COSMOS and previous methods [63, 88, 91] trained on Merged-30M datasets are compared in Tab. 4, alongside the OpenCLIP [13] models trained on larger datasets. These tasks are specifically designed to assess the perception capabilities of Vision-Language Models, as both visual and text input must be understood in detail to solve them.
Increasing the data size generally improves the performance of OpenCLIP models [13] as shown by the SugarCrepe results. Interestingly, COSMOS not only surpasses the previous baselines, but also outperforms the OpenCLIP models, demonstrating that self-distillation and cross-modality learning play crucial roles in the compositional understanding of multi-modal data. It is also noteworthy that COSMOS (trained on 30M samples) achieves the same performance (25.9%) as OpenCLIP [13] (trained on 1B samples) on the challenging benchmark MMVP-VLM [74]. This result suggests that COSMOS utilizes the pre-training data more efficiently, being able to learn distinguishing features with less data. Additionally, on the SugarCrepe dataset [33], where the task is to choose the correct caption among two captions (e.g., in the wrong caption some words are replaced, swapped inside the sentence or added), COSMOS achieves an average accuracy of 86.6%. This result signifies that COSMOS is able to differentiate between subtle differences in the text and, to a great extent, it does not treat the text as a bag-of-words. In Sec. E.6, we presents results in the MLLM setting by integrating our vision encoder in the LLaVA framework [49].
4.5 Ablation Study
Method ImageNet MSCOCO Flickr30K
Top-1 I2T T2I I2T T2I
1 CLIP 17.5 15.0 10.7 31.8 21.3
2 + Image Aug. 19.1 17.5 12.7 36.4 27.6
3 + EMA 19.1 17.5 12.7 36.5 27.6
4 + Image Self-distill. 22.8 20.6 14.6 40.2 29.4
5 + Text Aug. 34.4 50.4 37.5 81.3 65.5
6 + Text Self-distill. 35.4 51.0 38.0 82.5 67.1
7 + Cross-attn. (=COSMOS) 37.1 53.1 40.1 84.1 68.6
Table 5: Component ablation. We show top-1 accuracy on ImageNet [18], and R@1 score for I2T and T2I retrieval tasks on Flickr30K [84] and MSCOCO [48]. All models are trained on CC3M with the batch size of 1,024 and ViT-B/16 image encoder.
We validate the effectiveness of each component in Tab. 5, starting with a naive CLIP model trained on image and text pairs without augmentation (row 1). This simple setup achieves an accuracy of 15.0% and 10.7% in terms of zero-shot performance on MSCOCO image-to-text and text-to-image tasks. Applying image augmentation (row 2) improves all metrics, while EMA (row 3) shows marginal effects on the results. Applying self-distillation on the image encoder brings additional improvements, as it enforces local-to-global correspondence (row 4). Row 5 reveals that employing long synthetic captions as text augmentation plays an important role, boosting the accuracy on MSCOCO to 50.4% and 37.5%. Simply adapting self-distillation for the text encoder (row 6) slightly improves the performance (51.0% and 38.0% on MSCOCO). The self-distillation mechanism is further enhanced by applying a cross-attention module (row 7) reaching the best accuracy of 53.1% and 40.1% on MSCOCO retrieval task.
Additionally, we reproduce previous self-supervised methods based on CLIP, such as SILC [57] and SLIP [56], using the same synthetic dataset on which COSMOS is trained. We directly compare these methods with COSMOS reporting the results in Sec. E.1.
4.6 Qualitative Results
We present additional qualitative results of COSMOS in Fig. 3. We visualize the normalized attention weights of the cross-attention modules for both images and text. For clarity, we use only three colors for the captions, discretized by their original weight values (dark blue for low value, light green for intermediate value, and yellow for high value). Image attention results are based on the input text, while text attention results are conditioned on the input image.
The qualitative results illustrate that various objects are highlighted in both the image and the text, regardless of their location and size. In the top-left example, the sheep from the foreground and turbines from the background are accurately captured in both image and caption, despite occupying only a small part of the image. Additionally, our module can attend to specific attributes of objects, such as the wooden fence or the silver sports car in the bottom-right example. More qualitative results are in Sec. A.
Figure 3: Visualization of attention map in cross-attention modules. Attention weights are normalized between 0 and 1.
5 Conclusion
Our COSMOS, a model for vision-language pre-training, uses cross-modality representation for self-distillation. Specifically, we design a text-cropping strategy based on long caption datasets, which enable us to learn local-to-global text correspondences. Our cross-attention module for self-supervised learning further encourages our model to learn both visual and textual grounding, overcoming the previous shortcomings of the contrastive vision-language pre-training. Experimental results on various downstream tasks demonstrate that COSMOS outperforms strong baselines trained on much larger datasets on image-text retrieval, semantic segmentation and compositional understanding.
Acknowledgements. This work was partially funded by the ERC (853489 - DEXIM) and the Alfried Krupp von Bohlen und Halbach Foundation, which we thank for their generous support. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS [36] at Jülich Supercomputing Centre (JSC). We would also like to thank Shyamgopal Karthik and the entire EML group for helpful feedback and discussions.
Supplementary Material
Figure 4: Visualization of Attention Map. For different set of captions, we visualize the attention weights of the image and text cross-attention modules. The patch-wise (image) and token-wise (caption) attention weights are both normalized between 0 and 1.
In this supplementary material, we include detailed information about the training procedure of COSMOS. We first provide additional qualitative results in Sec. A. The dataset configurations are described in Sec. B, and the experimental setups are outlined in Sec. C. Additionally, we elaborate on the baseline configurations in Sec. D and present further experiments in Sec. E.
A Qualitative Results
In Fig. 4, we visualize the attention maps of the cross-attention modules with different captions to illustrate their focus areas. We first normalize the attention weights across all patches and map them back onto the original image. Similarly, tokens in the captions are colored based on the normalized attention weights. The qualitative results show that our cross-attention module enhances the learning of local representations in both modalities, namely image and text. For example, our model is capable of detecting relatively small objects in the image, such as sugar cubes (first example of the second row) or a broken smartphone (third example of the first row). This information is often overlooked by the previous models due to feature suppression.
B Dataset Configuration
As COSMOS is trained on web datasets, we had to download the images based on the corresponding URLs, but some of the URLs were invalid. Consequently, there is a discrepancy between the number of samples used to train our models and the original dataset sizes. The specific number of samples are reported in Tab. 6. We lost approximately 8.6% (YFCC15M) to 19.4% (CC12M) of samples compared to the original datasets due to expired URLs. This detail should be considered when directly comparing our models with others.
Dataset Original Ours Difference Percentage
CC3M [68] $,\!33$ 318 , 333 $2,\!823,\!019$ $500,\!229$ $85.1\%$
CC12M [9] $12,\!423,\!374$ $10,\!010,\!225$ $2,\!413,\!149$ $80.6\%$
YFCC15M [16] $15,\!388,\!848$ $14,\!065,\!827$ $1,\!323,\!021$ $91.4\%$
Merged-30M $31,\!130,\!555$ $26,\!899,\!071$ $4,\!236,\!399$ $86.4\%$
PixelProse [70] $16,\!896,\!423$ $15,\!037,\!386$ $1,\!859,\!037$ $89.0\%$
Table 6: Size of the pre-training datasets. We compare the original dataset sizes and the actual number of samples used to train our models. We also report the percentage of samples that we succeeded to retrieve from the original image URL links.
Additionally, COSMOS is evaluated on SugarCrepe [33], SVO [32], and MMVP [32], to assess the robustness of its multi-modal representations.
The SugarCrepe dataset [33] consists of images with positive and negative captions, where the model is required to choose the positive one given an input image (50% is the random chance). Negative captions are slightly different from the positive captions in terms of attributes and objects, with the goal of confusing the model. Since SugarCrepe is based on the COCO [48] validation set, we were able to download all images without any data loss.
SVO [32] consists of 36,841 data pairs, where each caption is paired with two images (one positive and one negative). Each image has its own triplet (subject, verb, object), and the positive and negative images differ in one component of the triplets. The original task is to match the caption with one of the two images based on the similarity scores given by VLMs (50% by random chance). Since naive CLIP [63] already achieves 80% accuracy, we increased the difficulty of this task by constructing negative captions. Based on the triplet of positive and negative images, we replaced the subject, verb, or object in the positive caption with the negative one. Currently, the model not only has to match the positive image with the positive caption but also match the negative image with the negative caption (25% by random chance). As SVO is released with image URL links, we were able to collect 22,220 data pairs (60.3% of the original size).
MMVP-VLM [32] contains 15 image-text pairs across 9 categories, resulting in a total of 135 image-text pairs. It categorizes 9 challenging visual patterns that most VLMs struggle with. Each data point consists of two images and two captions, requiring the VLMs to match the correct image and caption respectively (25% by random chance). Since Hendricks and Nematzadeh [32] released their dataset including images and captions, we were able to fully download the MMVP-VLM dataset as originally intended.
C Experiment Configuration
In this section, we detail the configuration of our experiments, including hyperparameters for training (Sec. C.1), the inference process for each downstream task (Sec. C.2), and the pseudocode of our training objective (Sec. C.3).
C.1 Hyperparameters
As shown in Tab. 7, we primarily follow the hyperparameters described in DreamLIP [91], with the exception of batch size. DreamLIP utilizes a batch size of 1,024 for CC3M and 8,192 for other datasets. Due to computational constraints, we adopt a batch size of 1,024 for CC3M and 4,096 for other datasets to train COSMOS. For reproducing CLIP [63] and SigLIP [88] with our setting, we use a batch size of 1,024 for CC3M and 6,144 for other datasets to establish a strong baseline.
Config Value
Optimizer AdamW [52]
Learning rate $5\times 10^{-4}$
Weight decay $0.5$
Adam $\beta$ $\beta_{1},\beta_{2}=(0.9,0.98)$
Adam $\epsilon$ $1\times 10^{-8}$
Total epochs $32$
Warm up iterations $2,000$
Learning rate schedule cosine decay
Table 7: Hyperparameter configuration.
As mentioned in the main paper, the teacher model is updated at each iteration using the exponential moving average (EMA) of the student model. To obtain an effective teacher for self-distillation, we need to determine the momentum parameter $\lambda$, which controls the update rule of the teacher parameter $\theta_{t}$ based on the student parameter $\theta_{s}$ (i.e., $\theta_{t}=\lambda\theta_{t}+(1-\lambda)\theta_{s}$). According to Caron et al. [7], a higher batch size requires a lower momentum parameter. Based on their configuration, we choose 0.999 for CC3M and 0.99 for other datasets by default. We did not explicitly perform a grid search for this parameter. Empirically, we found that fixing the momentum parameter obtains a better performance, unlike Caron et al. [7], who used a cosine scheduler for the momentum parameter which eventually converges to 1.0.
C.2 Inference
[…]
For zero-shot classification, we follow the process established by CLIP [63]. First, we construct prompts using the class label names for each dataset, as referenced in ALIP [82]. For each class, text embeddings are generated by the text encoder using different prompts, which are then averaged to obtain the final text embedding. Given an input image, the image encoder extracts the image embedding and calculates the cosine similarity scores between the image embedding and the text embeddings of all predefined classes in each dataset. The class label with the highest score is selected as the prediction.
Zero-shot retrieval is based on the image and text embeddings extracted from the pre-trained models, following the evaluation scheme presented in [82]. For each image-text pair, the pre-trained model generates image and text embeddings, collecting all embeddings for the entire dataset. We then compute a full cosine similarity matrix between all image embeddings and all text embeddings. Captions with the highest similarity scores are selected for each image to calculate the R@1, R@5, and R@10 metrics for image-to-text retrieval. Similarly, for text-to-image retrieval, we choose images with the highest scores for each caption. As each image in the validation or test set is equipped with multiple captions, image-to-text retrieval scores are generally higher than text-to-image scores.
In zero-shot semantic segmentation, we exclude the background category for PASCAL VOC [22] and PASCAL Context [55], following Cha et al. [8] and Wang et al. [77]. To be specific, we denote the original datasets with a background class as VOC21 and Context60, while the variants without the background category are referred to as VOC20 and Context59 in the main table. At inference time, for a given set of classes in the datasets, we obtain the corresponding text embeddings by querying our text encoder with a standard prompt. We compute the cosine similarity between the image patch embeddings (image tokens) and the text features of each class name to generate a segmentation map in a zero-shot manner. We adhere to the evaluation protocol established by SCLIP [77], including specifications for the window size, stride, and other parameters. We believe that the raw segmentation output of a VLM accurately reflects its zero-shot performance; therefore, we do not fine-tune our model or apply any post-refinement techniques such as PAMR [2].
To evaluate on SugarCrepe [33], SVO [32], and MMVP-VLM [74], we primarily referred to their evaluation demo prompts. Image and text embeddings are extracted from the pre-trained models, and then image-text pairs with higher cosine similarity scores are chosen as the final decision.
C.3 Pseudocode
To increase the clarity of our method, we present the pseudocode of the training objective in Algorithm 1. We empirically found that incorporating local image crops on the CLIP loss diminishes zero-shot performance, whereas using local text crops enhances it. Consequently, we compute the CLIP loss between all text crops and global image crops. We infer that integrating diverse captions during training allows the model to learn various objects shown in the images, while including local image crops during training likely leads to the misalignment of image-text pairs.
Algorithm 1 COSMOS: Pseudocode of our loss function
# img_g, img_l: Global&local crop of image
# txt_g, txt_l: Global&local crop of text
# Is, Ts: Student image&text encoder
# It, Tt: Teacher image&text encoder
# Generate embeddings of size [batch, seq_len, dim]
s_img_g, s_img_l = Is(img_g, img_l) # Student image
s_txt_g, s_txt_l = Ts(txt_g, txt_l) # Student text
t_img_g = It(img_g) # Teacher image
t_txt_g = Tt(txt_g) # Teacher text
# Split into ([CLS],img_tok) or ([EOT], txt_tok)
s_cls_g, s_img_tok_g = s_img_g
s_eot_g, s_txt_tok_g = s_txt_g
s_cls_l, _ = s_img_l
s_eot_l, _ = s_txt_l
t_cls_g, _ = t_img_g
t_eot_g, _ = t_txt_g
# Calculate CLIP loss
clip_loss = (sym_nce(s_cls_g, s_eot_g) +
sym_nce(s_cls_g, s_eot_l))/2
# Generate cross-modality embeddings
s_cls = {s_cls_g, s_cls_l} # Combine student crops
s_eot = {s_eot_g, s_eot_l} # Combine student crops
h_img = s_cls + cross-attn(q=s_cls, kv=s_txt_tok_g)
h_txt = s_eot + cross-attn(q=s_eot, kv=s_img_tok_g)
# Calculate cross-modality self-distillation loss
cosmos_loss = (sym_nce(h_img, t_cls_g) +
sym_nce(h_img, t_eot_g) +
sym_nce(h_txt, t_cls_g) +
sym_nce(h_txt, t_eot_g))/4
final_loss = clip_loss + cosmos_loss
  Notes: We assume one global view and one local view for simplicity. sym_nce represents the symmetric InfoNCE loss [60]. cross-attn denotes the cross attention module which requires key, value ($kv$) and query ($q$) as inputs.
D Baseline Configuration
In the main tables, we evaluate both the pre-trained weights from the official code repository and our reproduction. For DreamLIP [91], we used the official pre-trained weights with ViT/B-16 as the vision encoder. Since they did not provide weights for the ViT/B-32 vision encoder, we referenced the results from their table as shown in Tab. 10 and Tab. 11. For OpenCLIP [13], we utilized the models described in Tab. 8. Specifically, they trained their CLIP models on LAION-400M [66] until 12.8 billion examples were seen, using a batch size of 33,792 for both ViT-B/16 and ViT-B/32. For DataComp-1B [25], the models were trained until 12.8 billion examples were seen with a batch size of 90,112. Additionally, the models were trained on LAION-2B [67] with a batch size of 88,064 (ViT-B/16) or 79,104 (ViT-B/32) until 34 billion examples were seen.
Architecture Data Model Name
LAION-400M laion400m_e32
ViT-B/16 DataComp-1B datacomp_xl_s13b_b90k
LAION-2B laion2b_s34b_b88k
LAION-400M laion400m_e32
ViT-B/32 DataComp-1B datacomp_xl_s13b_b90k
LAION-2B laion2b_s34b_b79k
Table 8: OpenCLIP [13] model names.
E Extra Experiments
E.1 Comparison to Previous SSL Methods
Figure 5: Illustration of CLIP with self-supervised approaches. $I$ and $T$ denote the image and text encoders, respectively. $I_{t}$ (or $T_{t}$) and $I_{s}$ (or $T_{s}$) represent the teacher and student image (or text) encoders, where the teacher is an exponential moving average (EMA) of the student. (a) CLIP [63]: image and text embeddings are aligned during training. (b) SLIP [56]: contrastive loss is computed on sets of two different augmentations. (c) SILC [57]: self-distillation loss is obtained between local and global crops of the same image. (d) COSMOS: the cross-attention module is utilized to generate cross-modal representations which are optimized through the cross-modality self-distillation loss. We also design global and local crops of image and text pairs for effective self-supervised learning in VLMs.
Previous works, such as SLIP [56] or SILC [57], aim to enhance CLIP through self-supervision by explicitly optimizing local-to-global correspondences. These methods primarily focus on the image encoder. Fig. 5 illustrates how our approach differs from the self-supervised contrastive language-image pre-training methods, SLIP [56] and SILC [57].
In Tab. 9, we directly compare COSMOS with the previous methods depicted in Fig. 5. For a fair comparison, we re-implemented these methods based on the OpenCLIP [13] code repository and trained them on the CC3M dataset with long captions provided by DreamLIP [91]. The projection layer parameters follow the SLIP configuration, while the optimal temperature and loss scale for SILC were carefully determined. After a grid search, we selected student and teacher temperatures of 0.1 and 0.02, respectively, and used loss scale hyperparameters of (1.9, 0.1) for CLIP loss and self-distillation loss (see Tab. 20). Without extensive hyperparameter tuning, COSMOS significantly outperforms previous methods obtaining an accuracy of 37.1% in zero-shot classification on ImageNet [18], and 53.1% R@1 score in image-to-text retrieval on MSCOCO [48]. As mentioned in the introduction, focusing solely on enhancing image representation leads to sub-optimal results in VLMs, as shown by SILC which reached an accuracy score of only 30.4% on ImageNet and an R@1 score of 48.6% on image-to-text retrieval on MSCOCO, highlighting the importance of self-distillation with cross-modality representations.
Method ImageNet MSCOCO Flickr30k
Top-1 I2T@1 T2I@1 I2T@1 T2I@1
(a) CLIP [63] 23.9 40.2 27.2 68.4 52.1
(b) SLIP [56] 26.6 44.4 30.5 75.8 58.7
(c) SILC [57] 30.4 48.6 35.4 79.0 62.1
(d) COSMOS 37.1 53.1 40.1 84.1 68.6
Table 9: Comparison to methods described in Fig. 5. All models are trained on CC3M with long synthetic caption. We use the batch size of 1,024 and ViT-B/16 image encoder.
E.2 Experiments with ViT-B/32 Architecture
Flickr30K MSCOCO
Image-to-text Text-to-image Image-to-text Text-to-image
Data Method R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CC3M CLIP [63] 53.4 77.1 85.0 38.4 65.4 74.6 31.3 55.3 66.9 19.8 42.7 54.8
SigLIP [88] 56.4 78.9 85.7 40.2 66.5 76.0 30.8 55.5 66.8 20.6 43.9 55.4
DreamLIP [91] 57.6 84.4 89.6 42.2 69.0 77.7 33.4 60.7 72.0 23.4 47.2 58.6
COSMOS 74.3 92.3 95.9 59.2 82.6 89.1 47.6 73.1 82.0 33.5 59.7 70.6
CC12M CLIP [63] 76.8 92.9 95.8 60.3 83.4 89.4 50.6 76.5 84.3 34.9 61.5 72.3
SigLIP [88] 75.8 92.7 96.1 60.0 83.1 89.7 51.2 76.5 84.7 35.4 61.9 72.4
DreamLIP [91] 78.7 94.6 97.6 61.0 83.9 89.8 53.4 77.1 84.7 36.7 62.3 72.3
COSMOS 86.5 97.5 98.7 69.8 89.3 94.1 59.6 82.3 89.4 43.0 69.5 78.9
YFCC15M CLIP [63] 84.6 97.1 98.9 66.0 87.8 92.7 56.7 81.6 88.6 40.1 66.8 76.6
SigLIP [88] 82.5 97.4 98.5 66.9 87.8 92.3 56.3 82.0 89.0 40.0 66.7 76.9
DreamLIP [91] 84.9 97.3 98.7 66.0 86.4 91.4 55.7 80.5 88.2 39.8 66.0 75.5
COSMOS 90.2 98.7 99.4 73.3 91.6 95.4 64.5 86.1 91.8 46.0 72.2 81.0
Merged-30M CLIP [63] 85.6 96.7 99.0 69.5 89.9 94.3 59.3 83.1 89.9 42.8 69.3 79.0
SigLIP [88] 88.4 97.7 99.1 70.7 90.3 94.6 59.5 83.3 90.1 43.8 70.1 79.6
DreamLIP [91] 87.2 97.5 98.8 66.4 88.3 93.3 58.3 81.6 88.8 41.1 67.0 76.6
COSMOS 89.9 98.8 99.3 76.1 92.8 96.2 64.3 86.5 92.0 48.4 74.2 82.6
LAION-400M OpenCLIP [13] 79.7 95.0 97.6 60.9 84.8 90.7 51.7 76.6 84.9 33.7 59.4 69.9
DataComp-1B OpenCLIP [13] 80.1 94.6 97.2 62.9 85.4 91.1 54.6 78.4 85.8 36.3 62.1 72.6
LAION-2B OpenCLIP [13] 85.4 96.2 98.2 68.4 89.0 93.4 56.6 80.3 87.4 38.8 64.8 74.7
Table 10: Zero-shot image-text retrieval results in terms of R@1, R@5, and R@10 on the Flickr30K [84] and MSCOCO [48] datasets. The vision encoder architecture is ViT-B/32. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
In addition to the experiments with ViT-B/16 presented in the main paper, we conducted the same experiments using the ViT-B/32 vision encoder architecture, as shown in Tab. 10 and Tab. 11. Overall, the improvements achieved with COSMOS are consistent with those observed in the main table using ViT-B/16.
In the zero-shot retrieval tasks (Tab. 10), COSMOS not only surpasses previous strong baselines (CLIP [63], SigLIP [88], and DreamLIP [91]) trained on the same datasets, but also exceeds the performance of OpenCLIP [13] models trained on much larger datasets. Notably, COSMOS trained on Merged-30M achieves 64.3% and 48.4% in image-to-text and text-to-image R@1 scores on MSCOCO, significantly outperforming DreamLIP (58.3% and 41.1%). Furthermore, COSMOS trained on CC12M already demonstrates higher accuracy compared to OpenCLIP trained on LAION-2B dataset. These results highlight the effectiveness of COSMOS in generating more fine-grained and comprehensive multi-modal representations through cross-modality self-distillation.
In the zero-shot classification tasks (Tab. 11), COSMOS improves all metrics compared to CLIP [63] and SigLIP [88] across most datasets, while achieving comparable results to DreamLIP [91] on YFCC15M and Merged-30M. DreamLIP [91] aligns visual patch embeddings with positive text embeddings, promoting the learning of global structure by image tokens, which enhances the performance of classification task where the model focuses on global information. Although COSMOS optimizes local-to-global correspondence to enhance the representation of local information through self-distillation, its performance is similar to DreamLIP. Additionally, increasing the dataset size generally improves accuracy, suggesting that the total number of images is a crucial factor for achieving high performance in classification. The results obtained with ViT-B/32 are consistent with those of ViT-B/16 from the main paper.
E.3 Experiment on PixelProse Dataset
To demonstrate the adaptability of COSMOS to various captions, we employ PixelProse [70], a synthetic caption dataset similar to DreamLIP. PixelProse filters and combines data from CommonPool [25], CC12M [68], and RedCaps [19] to create over 16 million image and alt-text pairs. They used Gemini 1.0 Pro [70] to generate new captions. As shown in Tab. 12 and Tab. 13, we trained CLIP [63], SigLIP [88], and COSMOS on PixelProse using the same setup as before. The actual number of image-text pairs used to train these models is detailed in Tab. 6.
In Tab. 12, COSMOS consistently outperforms the CLIP and SigLIP models by a significant margin with both ViT-B/16 and ViT-B/32 vision encoders. For instance, COSMOS with ViT-B/16 achieves 62.4% and 43.4% in R@1 scores for image-to-text and text-to-image retrieval on MSCOCO. This performance even surpasses strong baselines from OpenCLIP [13], which achieve 56.5% and 37.9% on LAION-400M, 58.2% and 39.8% on DataComp-1B, and 59.3% and 41.7% on LAION-2B (see Tab. 1 main).
Similarly, Tab. 13 shows that COSMOS outperforms previous baselines in the zero-shot classification tasks. Notably, COSMOS trained on PixelProse achieves the best average accuracy of 60.7% with ViT-B/16 and 58.0% with ViT-B/32 among all models trained on the synthetic caption datasets (e.g., CC3M, CC12M, YFCC15M, and Merged-30M). However, the same models trained on PixelProse do not achieve the best accuracy in retrieval tasks according to Tab. 1 main and Tab. 10. We infer that this might be due to the length and content of the captions used during the training. DreamLIP utilizes three MLLMs to generate long synthetic captions, resulting in longer captions than those generated by PixelProse, which only uses Gemini Pro. Consequently, models trained on longer and more varied captions are better at capturing local information (i.e., better at retrieval task), while models trained on relatively shorter and similar captions are better at focusing on global objects in the image (i.e., better at classification task). Investigating this trade-off based on the nature of these long captions would be an intriguing direction for future work.
Data Model
Food-101
CIFAR-10
CIFAR-100
SUN397
Cars
Aircraft
DTD
Pets
Caltech-101
Flowers
ImageNet
Average
CC3M CLIP [63] 14.2 73.5 37.8 39.3 2.4 1.4 14.4 16.9 66.4 10.7 19.7 27.0
SigLIP [88] 13.5 74.2 40.1 39.8 2.8 1.3 12.9 16.3 67.3 11.3 21.2 27.3
DreamLIP [91] 16.1 82.0 45.4 41.3 2.5 1.0 13.9 18.8 64.4 14.1 25.9 29.6
COSMOS 24.9 80.0 51.9 51.5 3.8 1.5 25.2 29.9 77.2 21.9 33.0 36.4
CC12M CLIP [63] 41.6 88.5 57.5 54.5 11.3 2.3 23.4 43.3 80.4 15.7 37.4 41.4
SigLIP [88] 41.9 87.5 59.3 55.2 11.8 1.6 27.3 41.8 80.9 18.4 37.8 42.1
DreamLIP [91] 48.9 86.4 63.0 55.7 17.9 1.9 23.5 41.9 83.2 25.8 44.2 44.8
COSMOS 52.9 91.2 67.5 61.0 23.8 3.7 32.1 54.2 85.5 30.6 46.7 49.9
YFCC15M CLIP [63] 38.9 86.2 58.2 53.3 7.1 4.0 23.9 27.6 76.8 38.0 38.9 41.2
SigLIP [88] 37.7 86.1 57.1 53.2 6.4 4.3 25.3 30.4 77.4 35.3 38.6 41.1
DreamLIP [91] 51.7 87.9 60.7 54.8 9.4 7.1 26.8 36.3 79.6 48.6 46.6 46.3
COSMOS 40.3 84.9 57.0 54.6 13.5 5.9 31.3 38.6 82.1 47.8 48.1 45.8
Merged-30M CLIP [63] 54.8 90.0 67.1 62.0 13.0 3.6 27.6 49.4 83.8 41.4 45.6 48.9
SigLIP [88] 52.8 90.8 66.1 63.4 15.0 5.4 29.9 47.8 84.4 35.7 46.5 48.9
DreamLIP [91] 68.2 91.8 69.2 62.2 20.7 8.0 32.1 62.8 86.1 48.5 55.7 55.0
COSMOS 65.9 91.5 70.8 64.6 23.4 7.6 37.6 57.3 86.1 52.2 53.4 55.5
LAION-400M OpenCLIP [13] 78.2 88.4 68.3 65.0 74.5 14.6 52.4 84.9 88.4 65.9 60.2 67.3
DataComp-1B OpenCLIP [13] 86.3 95.6 80.4 67.3 87.3 24.8 57.2 90.2 91.6 73.2 69.2 74.8
LAION-2B OpenCLIP [13] 82.7 93.6 75.8 68.7 86.1 24.5 55.8 90.4 90.5 71.6 66.5 73.3
2.5B MetaCLIP [80] 82.7 95.2 77.7 66.8 77.4 27.0 58.9 90.9 92.8 69.9 67.6 73.4
Llip [41] 84.1 95.5 80.8 68.6 82.2 34.9 58.8 92.3 92.9 74.8 67.5 75.6
Table 11: Zero-shot classification results in terms of top-1 accuracy on the ImageNet [18] and Food101 [4], CIFAR-10 [39], CIFAR-100 [39], SUN397 [79], Stanford Cars [38], FGVCAircraft [54], DTD [14], Oxford Pets [61], Caltech101 [24], Flowers102 [58], and ImageNet [18] datasets. The vision encoder architecture is ViT-B/32. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
Flickr30K MSCOCO
Image-to-text Text-to-image Image-to-text Text-to-image
Method R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Model Architecture: ViT-B/32
CLIP [63] 76.1 92.1 95.7 57.5 81.8 88.6 49.4 73.7 82.2 32.0 57.1 68.2
SigLIP [88] 74.4 92.2 96.1 56.8 81.3 87.7 48.9 74.4 82.2 32.0 57.2 68.0
COSMOS 85.6 96.9 98.3 66.3 87.6 92.7 57.2 80.9 88.0 38.9 64.8 74.6
Model Architecture: ViT-B/16
CLIP [63] 85.2 96.2 98.6 66.3 87.8 93.2 56.9 79.7 87.4 38.0 64.1 74.5
SigLIP [88] 85.4 96.7 97.8 66.6 88.4 93.5 57.5 80.5 87.6 38.9 64.5 74.4
COSMOS 89.9 98.5 99.5 73.6 92.0 95.4 62.4 84.0 89.7 43.4 69.3 78.8
Table 12: PixelProse: Zero-shot image-text retrieval results in terms of R@1, R@5, and R@10 on the Flickr30K [84] and MSCOCO [48] datasets. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison.
Model
Food-101
CIFAR-10
CIFAR-100
SUN397
Cars
Aircraft
DTD
Pets
Caltech-101
Flowers
ImageNet
Average
Model Architecture: ViT-B/32
CLIP [63] 52.4 85.2 56.1 54.1 26.0 1.9 32.0 59.0 85.0 34.4 44.9 48.3
SigLIP [88] 53.8 85.7 59.4 55.6 25.2 3.1 32.1 64.1 84.8 37.1 44.6 49.6
COSMOS 66.3 89.9 67.4 60.4 47.0 4.4 37.3 76.6 89.5 45.1 54.3 58.0
Model Architecture: ViT-B/16
CLIP [63] 63.2 87.9 60.2 59.1 37.0 2.3 40.1 68.8 86.6 42.1 50.8 54.4
SigLIP [88] 63.4 87.5 58.6 60.5 36.2 3.5 37.8 63.5 87.0 43.4 51.1 53.9
COSMOS 72.1 92.1 71.9 60.8 53.0 3.6 43.4 74.4 90.1 46.8 59.6 60.7
Table 13: PixelProse: Zero-shot classification results in terms of top-1 accuracy on ImageNet [18] and 10 common downstream datasets. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison.
Method ImageNet MSCOCO Flickr30k Training
Top-1 I2T@1 T2I@1 I2T@1 T2I@1 Time Mem.
1 CLIP: 1 globals + 0 locals 24.5 38.9 26.8 69.1 51.2 3.8h 10.5G
2 CLIP: 2 globals + 0 locals 29.7 47.6 33.8 78.1 60.7 12.2h 16.3G
3 CLIP: 2 globals + 2 locals 31.3 49.6 35.9 78.4 62.1 17.6h 20.6G
4 COSMOS: 2 globals + 0 locals 31.0 50.1 35.4 80.9 63.2 14.6h 18.2G
5 COSMOS: 2 globals + 2 locals 34.8 51.5 38.5 82.7 65.9 21.4h 23.4G
6 COSMOS: 2 globals + 4 locals 36.3 52.3 39.6 83.5 67.4 26.5h 28.2G
7 COSMOS: 2 globals + 6 locals 37.1 53.1 40.1 84.1 68.6 31.7h 32.6G
Table 14: Ablation on the training efficiency of COSMOS. Total running time and peak memory usage per GPU are shown with different configuration of global and locap crops. We also report zero-shot classification results on ImageNet [18] and zero-shot retrieval results on Flickr30K [84] and MSCOCO [48]. We train our model (ViT-B/16) on four 4-GPU machines using the CC3M dataset with the batch size of 1,024.
E.4 Ablation Study on Training Efficiency
In Tab. 14, we report the GPU memory requirements and training time based on the number of global and local crops, along with the corresponding zero-shot performance. Compared to CLIP, COSMOS requires 10-20% more training time and GPU memory due to its teacher-student setup (row 2 vs row 4, and row 3 vs row 5). As no gradients flow into the teacher during training, COSMOS does not add too much computational overhead while outperforming CLIP.
For COSMOS, we fixed the number of global crops for both image and text to two, while varying the number of local crops included during training (row 4-row 7). As a result, increasing the number of local crops improves zero-shot classification and retrieval tasks, achieving 37.1% accuracy on ImageNet and 53.1% and 40.1% R@1 scores on image-to-text and text-to-image retrieval on the MSCOCO validation set with six local crops. The improvement is most significant between zero and two local crops (row 4 vs row 5), with diminishing returns as more local crops are added. GPU memory usage and training time also increase with the number of local crops. Therefore, one could determine the optimal number of local crops based on the available computational resources.
E.5 Experiment on the Winoground Dataset
Method Data Size Text Image Group
CLIP [63] 30M 29.3 13.0 8.5
SigLIP [88] 30M 29.0 11.0 8.3
DreamLIP [91] 30M 27.8 15.8 11.5
COSMOS 30M 30.8 16.5 11.3
400M 25.5 11.5 7.8
OpenCLIP [13] 1B 29.8 8.8 7.3
2B 28.0 10.8 8.3
Table 15: Evaluation on Winoground [73].
In addition to the visual perception and contextual understanding tasks presented in the main paper, we also evaluate COSMOS on the Winoground dataset [73]. Each entry in the dataset consists of two images and two captions, where the task is to correctly match the image-text pairs (25% by random chance). Both captions contain an identical set of words, but in different orders, assessing whether the model possesses sufficient compositional understanding for challenging image-text pairs.
In Tab. 15, we compare COSMOS with three baselines (CLIP [63], SigLIP [88], and DreamLIP [91]) as well as OpenCLIP [13] models trained on larger datasets. The results of the OpenCLIP models indicate that scaling up the size of pre-training datasets does not necessarily improve the performance. In other words, large datasets alone cannot ensure high compositional knowledge in VLMs. However, COSMOS outperforms all OpenCLIP models, suggesting that our self-distillation and cross-modality learning effectively enhance the model’s contextual understanding. COSMOS achieves 30.8% and 16.5% in text and image scores, respectively, surpassing the previous best baselines of 29.3% text score from CLIP and 15.8% image score from DreamLIP, while achieving comparable results in terms of group score.
E.6 Experiment on MLLM setting
Method Data Size ScienceQA POPE GQA TextVQA MMMU
CLIP [63] 30M 65.2 80.1 58.7 53.5 33.9
OpenCLIP [13] 400M 67.2 81.3 59.8 54.4 36.5
COSMOS 30M 67.8 83.2 60.4 55.3 36.8
Table 16: LLaVA v1.5 experiment. CLIP and COSMOS trained on Merged-30M and OpenCLIP trained on LAION-400M.
In order to evaluate COSMOS on MLLM setting, we adapt our vision encoder (ViT-B/16) to LLaVA framework [49]. We follow the training process of LLaVA v1.5 which consists of feature alignment pre-training and visual instruction tuning. Various benchmarks are selected for comprehensive evaluation including ScienceQA [53], POPE [46], GQA [34], TextVQA [69], and MMMU [86]. In Tab. 16, COSMOS outperforms both CLIP trained on the same data (Merged-30M) and OpenCLIP trained on LAION-400M, which demonstrates the effectiveness of COSMOS on visual reasoning and compositional question answering.
E.7 Ablation Study on Text Cropping Strategy
Method ImageNet MSCOCO Flickr30K
Top-1 I2T@1 T2I@1 I2T@1 T2I@1
Masked text [45] 28.5 46.0 32.8 76.2 59.2
Summarized text [26] 33.9 51.7 37.8 81.0 65.3
Local within global 34.6 52.1 37.8 81.2 66.1
COSMOS 35.1 52.6 38.9 83.0 66.5
Table 17: Ablation on text cropping strategies. We compare various text cropping methods in terms of zero-shot classification results on ImageNet [18] and zero-shot retrieval results on Flickr30K [84] and MSCOCO [48]. All models are trained on CC3M with batch size 1,024 and ViT-B/16 image encoder.
In our paper, both global and local crops of captions are randomly sampled from long synthetic captions. Typically, global texts (1-5 sentences) are longer than local texts (1 sentence). The sampling processes of global and local crops are entirely independent. Therefore, global captions may or may not include local captions, similar to the image crop method in SimCLR [11] and DINO [7]. While mismatches can occur between global and local captions via random sampling, as with the global and local crops of images, the model can learn conceptual correspondences. For example, for an image of a park, the global text may describe the park, while the local text describes a dog. The model then learns that "park" and "dog" are related in the text.
[…]
E.8 Additional Results on Semantic Segmentation
Method VOC20 City. Context59 ADE COCO-Stf.
CLIP [63] 11.3 5.0 4.5 1.3 2.8
SigLIP [88] 14.5 5.5 5.8 2.2 3.8
DreamLIP [91] 1.8 0.9 0.4 0.1 0.1
COSMOS 53.6 13.9 15.7 8.5 10.7
Table 18: Zero-shot semantic segmentation results in terms of mean Intersection over Union (mIoU). The vision encoder architecture is ViT-B/16 and all models are trained on Merged-30M.
In addition to Table 3 in main paper, we compare segmentation performance of COSMOS to CLIP [63], SigLIP [88], and DreamLIP [91] in Tab. 18. Surprisingly, DreamLIP performs poorly in every semantic segmentation benchmarks, likely because its loss function weakens local image representation by matching local visual patches with global text. Ours is by far the best performing method, likely due to its cross-modality embedding and local-to-global matching, alleviating local feature suppression.
E.9 Additional SOTA Comparison
Method Data Size Batch Size MSCOCO Flickr30K
I2T T2I I2T T2I
VeCLIP [40] 300M 32k 67.8 48.9 91.2 76.3
MobileCLIP-B [76] 1B 65k 68.8 50.6 91.4 77.3
COSMOS 30M 4k 68.0 52.5 92.9 80.3
Table 19: Comparison to VeCLIP [40] and MobileCLIP [76] in terms of zero-shot retrieval results on Flickr30K [84] and MSCOCO [48].
In Table 1 and 2 in the main paper, COSMOS is compared to other methods using multi-modal data augmentation including DreamLIP [91], LaCLIP [23], and MLLM-A [51]. We additionally report the results of VeCLIP [40] and MobileCLIP [76] for comparison. VeCLIP exploits LLaVA [50] to generate detailed captions while MobileCLIP utilizes CoCa [85] to generate multiple synthetic cations. Although COSMOS is trained with much smaller pre-training set and batch size, it outperforms other methods, demonstrating the efficient usage of synthetic captions within our framework.
E.10 Rescaling Loss in Previous Methods
Method ImageNet MSCOCO Flickr30k
Top-1 I2T@1 T2I@1 I2T@1 T2I@1
CLIP w/ Aug. EMA 20.0 19.2 14.7 38.7 29.3
SILC (1.0, 1.0) 14.7 9.2 13.4 26.8 18.7
SILC (1.5, 0.5) 19.0 17.4 12.9 33.8 25.2
SILC (1.8, 0.2) 21.0 20.8 14.8 40.2 29.3
SILC (1.9, 0.1) 21.4 21.1 15.3 42.0 29.7
SILC (1.95, 0.05) 21.4 20.4 15.1 40.2 29.5
Table 20: Ablation on the loss scale of SILC [57]. SILC with different loss scale (a,b) where the total loss is calculated as $\mathcal{L}_{\text{total}}=a\mathcal{L}_{\text{CLIP}}+b\mathcal{L}_{\text{self-distill}}$. Models are trained on CC3M with the batch size of 1,024 and one global and one local crops are used as augmentation.
In Tab. 20, we conduct an experiment to demonstrate the effect of loss scaling in SILC [57], which also utilizes self-supervision in contrastive vision-language pre-training. We adjust the scale of the CLIP loss ($a$) and the self-distillation loss ($b$), while maintaining their sum constant (i.e., $a+b=2$). For comparison, we also include CLIP [63] with the same augmentation and EMA. Interestingly, a naive summation of the two losses (i.e., $a=1.0,b=1.0$) results in a worse performance compared to CLIP, highlighting the importance of selecting appropriate scaling parameters for optimal performance. Consequently, $a=1.9$ and $b=0.1$ yield the best results, which we used to reproduce their results in Tab. 9.
Similarly, other works [45, 56, 21, 65] that integrated self-supervised learning in VLM training adopt loss scaling parameters to balance the learning speed between the contrastive objective and the self-distillation objective. This is due to the different scales of the loss functions, as previous works often employ the symmetric InfoNCE loss [60] for contrastive learning while using the cross-entropy loss between masked inputs or global and local crops for self-supervised learning. We unify the loss function with the InfoNCE loss, eliminating the need for a scaling factor, since the CLIP loss and the COSMOS loss are already updated on the same scale.